# Part II: Variational bootstrapping of the $\Lambda$CDM model

### v1

@@ V.5 Analysis of the equations
Before closing our discussion, let us comment on the level of generality of the obtained expressions and on how these relates to previously formulated models in metric-affine gravity. Because of the freedom of choice in $\Lambda$ and $\vec{V}$, our Lagrangian (46) still encompasses many of the models present in the literature. For a state of the art on metric-affine theories of gravity, we redirect the reader to [9] while, for a review on historical developments, one can see [7, 14].
@@
the choise
$\Lambda(\boldsymbol{g},\mathbf{L},\accentset{\circ}{\nabla}\mathbf{L})=-\frac{1}{2}\left(L^{\mu}_{\phantom{\mu}\nu\left[\mu\right|}L^{\nu\rho}_{\phantom{\nu\rho}\left|\rho\right]}+\accentset{\circ}{\nabla}_{\left[\mu\right|}L^{\mu\nu}_{\phantom{\mu\nu}\left|\nu\right]}\right),$ (72)
@@
where each term is $(-1)$-homogeneous with respect to $\boldsymbol{g}$, and $\vec{V}(\boldsymbol{g},\mathbf{L},\accentset{\circ}{\nabla}\mathbf{L})=\vec{0}$ leads to the desired Lagrangian.
@@
A more interesting example is that of quadratic metric-affine theories [17, 18], i.e. theories whose Lagrangian contains terms quadratic in the curvature $\mathbf{R}$, torsion $\mathbf{T}$ and/or non-metricity $\mathbf{Q}$ of $\Gamma$. In this case, one can easily see that appropriate choices of $\Lambda$ and $\vec{V}$ can generate all the terms under consideration in [17] or [18] except those who are quadratic in $\mathbf{R}$e.g. the term proportional to the coefficient $z_{4}$ in equation (3.1) of [17]. This comes from our hypothesis 1 about the form of $\Lambda$ and $\vec{V}$ and the fact that our Lagrangian contains only a term linear in $\accentset{\circ}{R}$.
@@
Following the same line, our procedure then also selects a restrictive class of quadratic gauge theories of gravity within a metric-affine framework as it constrain the dependence in $\mathbf{R}$ (Compare with chapter 9 of [7]).
@@
Finally, let us mention that, while we here considered theories based on a fully general metric-affine setup, many models studied in the literature start with some restriction on the connection $\Gamma$; typically $\mathbf{Q}\equiv\mathbf{0}$, as in Einstein-Cartan gravity or Poincaré gauge gravity, see respectively chapters 4 and 5 of [7]. Naturally, this would not affect the behavior of our procedure and the conclusion would remain the same [^10] Technically, imposing restrictions on the connection at the kinematical level account for changing the configuration manifold on which our Lagrangian theory is based. For example, by imposing $\mathbf{Q}\equiv\mathbf{0}$ at the kinematical level, one assumes the configuration manifold to be $Y=\operatorname{Met}(M)\times_{M}\operatorname{Tor}(M),$ where $\operatorname{Tor}(M)$ denotes the bundle of torsion – i.e. the bunble of type $(1,2)$ tensor fields antisymmetric over their two lower indices – instead of $Y=\operatorname{Met}(M)\times_{M}T^{1}_{2}M$ as in this paper, see [12]. Nevertheless, since we are interested in variational bootstrapping with respect to the metric anyway, the computation and conclusions would perform exactly the same as in our case.: Variational bootstrapping of the $\Lambda$CDM equations is compatible with quadratic metric-affine theories of gravity except for terms that are non-linear in the curvature tensor $\mathbf{R}$.
@@
As a final comment, let us emphasize that the particularity of the method outlined here in this business of finding Lagrangians is that it directly relates the notion of “minimality” to the variational structure of the theory. In other words, variational bootstrapping allows us to select metric affine models which are minimal extensions of the $\Lambda$CDM model in a variational sense and this, as we just saw, allows for a more restrictive selection of Lagrangians.
@@ VI Conclusion
In this article, we studied the application of variational bootstrapping, first discussed in [12], to metric-affine theories of gravity. Through a carreful examination of the $\Lambda$CDM model of cosmology, we proposed a minimal modification of GR’s field equation, based on that model, in the context of a metric-affine theory. Our hypotheses for that construction are explicited in section IV (see also section V.2). Then, variational bootstrapping allowed us to construct a Lagrangian for the full theory, leading to “corrected” metric equations (66) but also giving access to field equations for the distortion tensor (independent connection) (70).
We showed in section V.5 how our model related to previously formulated ones. We saw that the notion of “variationaly minimal extension” outlined in this paper gives a more restrictive criterion to select Lagrangians than the ones usually consiered in the literature when considering e.g. quadratic metric-affine theories. Naturally, this does not rule out the other existing models per se ! It should nevertheless be regarded as an interesting property since, in modified gravity in general (and in metric-affine gauge theories of gravity in particular), one usually struggles to single out a theory exihibiting a desired behaviour.
This investigation demonstrates that variational bootstrapping is a powerful tool to construct theoretical models based on phenomenological requirements. The most remarkable feature of the method is that it allows to obtain a sensible Lagrangain for the theory under study by only guessing a part of the system’s dynamics.
An important direction to extend the present work is to investigate solutions to the field equations (66$\land$(70). In particular, it would be interesting to see if a well-chosen form of the scalar function $\Lambda$ and vector field $V$ would allow, in a cosmological setup, to dynamically produce a behavior of the scale factor $a(t)$ compatible with current universe’s acceleration and/or with inflation.
In addition to that, it would also be necessary to study the behavior of compact objects (black holes, neutron stars) in our theory to see how these differ from GR’s ones and to which extend this remains compatible with modern observations.
@@ A.1 Symmetry
$\Phi:G\times M\to M:(g,x)\mapsto\Phi(g,x)\eqqcolon\Phi_{g}(x)\eqqcolon\Phi^{x}(g),$
@@
When $G$ is a Lie group, it is usually more convenient to study this notion of symmetry in terms of the infinitesimal action of the symmetry group. Let us denote by $\mathfrak{g}$ the Lie algebra of $G$ and by $X_{\xi}$ the generating vector field for a given $\xi\in\mathfrak{g}$111111i.e. $X_{\xi}$ is the vector field whose integral curves are given by the flow of $\Phi$ in the sense that, given a point $p\in M$, the integral curve of $X_{\xi}$ through $p$ is $\mathscr{C}^{p}_{X_{\xi}}:t\mapsto\Phi(\exp(\xi t),p),$ where $\exp$ is the exponential map., if a tensor $\boldsymbol{T}$ is symmetric in the sense of (73) then
@@
If $G$ is connected, conditions (73) and (74) are equivalent and one can then derive the conditions for a tensor field $\boldsymbol{T}$ to present a given symmetry via (74). It is the condition that we will use in the following.
Following this definition of symmetry, a metric-affine spacetime will be symmetric under a given group $G$ provided both the metric $\boldsymbol{g}$ and the distortion $\mathbf{L}$ are invariant under the action of $G$.
@@ A.2 Generators of cosmological symmetry
Cosmological symmetry refers to the idea that, following the cosmological principle, no observer has a prefered place in the universe so that spacetime must be spatially homogeneous and isotropic. More precisely here, we then expect our setup to be invariant under both the “spatial” rotation group $SO(3)$ (providing isotropy, at least around one point) and a group representing “spatial translations” (ensuring isotropy around actually any point and hence homogeneity). One thus obtains 6 generators of cosmological symmetry $X_{i}$ $(i=1,\cdots,6)$. In local spherical coordinates $\left(x^{\mu}\right)=\left(t,r,\theta,\varphi\right)$ on $M$, $SO(3)$-generators are described as:
@@
as they satisfy the commutation relation of $\mathfrak{so}(3)$ :
$\left[X_{i},X_{j}\right]=\epsilon_{ijk}X_{k},$ (76)
[…]
Generators of (quasi-)translations are:
@@
where $k\in\Set{-1,0,1}$ is the sign of the spatial curvature. When $k=0$ (flat spatial geometry), $X_{1}$, $X_{2}$ and $X_{3}$ correspond to the expression, in spherical coordinates, of the usual generators of spatial rotation around respectivelly the $x$-, $y$- and $z$-axis of an orthonornmal cartesian frame, while $X_{4}$, $X_{5}$ and $X_{6}$ correspond to the usual generators of spatial translation along respectivelly the $x$-, $y$- and $z$-axis.
@@ B.2.1 Variations with respect to the metric $\boldsymbol{g}$
When computing variation with respect to the metric, one has
@@
$\delta_{g}\accentset{\circ}{\Gamma}^{\rho}_{\mu\nu}$ $=\frac{1}{2}g^{\rho\alpha}\left(\accentset{\circ}{\nabla}_{\nu}\delta g_{\alpha\mu}+\accentset{\circ}{\nabla}_{\mu}\delta g_{\nu\alpha}-\accentset{\circ}{\nabla}_{\alpha}\delta g_{\mu\nu}\right)$ (92)
$=g^{\rho\alpha}\ \accentset{\circ}{\nabla}_{\lceil\mu}\delta g_{\nu\alpha\rfloor},$ (93)

### v2

@@ V.5 Analysis of the equations
Before closing our discussion, let us comment on the level of generality of the obtained expressions and on how these relates to previously formulated models in metric-affine gravity. Because of the freedom of choice in $\Lambda$ and $V$, our Lagrangian (46) still encompasses many of the models present in the literature. For a state of the art on metric-affine theories of gravity, we redirect the reader to [9] while, for a review on historical developments, one can see [7, 14].
@@
the choice131313where each term is $(-1)$-homogeneous with respect to $\boldsymbol{g}$
$\Lambda(\boldsymbol{g},\mathbf{L},\accentset{\circ}{\nabla}\mathbf{L})=-\frac{1}{2}\left(L^{\mu}_{\phantom{\mu}\nu\left[\mu\right|}L^{\nu\rho}_{\phantom{\nu\rho}\left|\rho\right]}+\accentset{\circ}{\nabla}_{\left[\mu\right|}L^{\mu\nu}_{\phantom{\mu\nu}\left|\nu\right]}\right)$ (73)
and $V(\boldsymbol{g},\mathbf{L},\accentset{\circ}{\nabla}\mathbf{L})=0$ in the seed equation (39) will lead, after variational bootstrapping, to a Lagrangian given by the right-hand side of (72).
@@
A more interesting example is that of quadratic metric-affine theories [17, 18], i.e. theories whose Lagrangian contains terms that are quadratic in the curvature $\mathbf{R}$ and/or torsion $\mathbf{T}$ and/or non-metricity $\mathbf{Q}$ of $\Gamma$. In this case, one can easily see that appropriate choices of $\Lambda$ and $V$ can generate all the terms under consideration in [17] or [18] except those which are quadratic in $\mathbf{R}$ (e.g. the term proportional to the coefficient $z_{4}$ in equation (3.1) of [17]). This comes from our hypothesis 1 on the form of $\Lambda$ and $V$, which implies that our Lagrangians can only contain a linear term in $\accentset{\circ}{R}$.
@@
Following the same line, our procedure then also selects a restrictive class of quadratic gauge theories of gravity within a metric-affine framework as it constrains the dependence in $\mathbf{R}$ (Compare with chapter 9 of [7]).
@@
Finally, let us mention that, while we here considered theories based on a fully general metric-affine setup, many models studied in the literature start with some restriction on the connection $\Gamma$; typically $\mathbf{Q}\equiv\mathbf{0}$, as in Einstein-Cartan gravity or Poincaré gauge gravity, see respectively chapters 4 and 5 of [7]. Naturally, this would not affect the behavior of our procedure and the conclusion would remain the same [^14] Technically, imposing restrictions on the connection at the kinematical level account for changing the configuration manifold on which our Lagrangian theory is based. For example, by imposing $\mathbf{Q}\equiv\mathbf{0}$ at the kinematical level, one assumes the configuration manifold to be $Y=\operatorname{Met}(M)\times_{M}\operatorname{Tor}(M),$ where $\operatorname{Tor}(M)$ denotes the bundle of torsions – i.e. the bundle of type $(1,2)$ tensor fields antisymmetric over their two lower indices – instead of $Y=\operatorname{Met}(M)\times_{M}T^{1}_{2}M$ as in this paper, see [12]. Nevertheless, since we are interested in variational bootstrapping with respect to the metric anyway, the computation and conclusions would perform exactly the same as in our case.: Variational bootstrapping of the $\Lambda$CDM equations is compatible with quadratic metric-affine theories of gravity except for terms that are non-linear in the curvature tensor $\mathbf{R}$.
@@
As a final comment, let us emphasize that the particularity of the method outlined here in this business of finding Lagrangians is that it directly relates the notion of “minimality” to the variational structure of the theory. In other words, variational bootstrapping allows us to select metric-affine models which are minimal extensions of the $\Lambda$CDM model in a variational sense and this, as we just saw, allows for a more restrictive selection of Lagrangians.
@@ VI Conclusion
In this article, we studied the application of variational bootstrapping, first discussed in [12], to metric-affine theories of gravity. Through a careful examination of the $\Lambda$CDM model of cosmology, we proposed a minimal modification of GR’s field equation, based on metric-affine geometry; our hypotheses for that construction are explained in section IV (see also subsection V.2). Then, variational bootstrapping allowed us to systematically construct the Lagrangian for the full theory, leading to the “corrected” metric equations (66) and, at the same time, completely determining the field equations for the distortion tensor (independent connection) (71).
We showed in subsection V.5 how our model (which is actually, a whole family of metric-affine models, obeying certain minimality criteria) relates to previously formulated ones. We saw that the notion of “variationaly minimal extension” outlined in this paper gives a more restrictive criterion to select Lagrangians than the ones usually considered in the literature when considering e.g., quadratic metric-affine theories. Naturally, this does not rule out the other existing models per se! It should nevertheless be regarded as an interesting property since, in modified gravity in general (and in metric-affine gauge theories of gravity in particular), one usually struggles to single out a theory exhibiting a desired behavior.
Last but not least, this investigation demonstrates that variational bootstrapping is a powerful tool to construct theoretical models based on phenomenological requirements. The most remarkable feature of the method is that it allows to obtain a sensible Lagrangian for the theory under study by only guessing a part of the system’s dynamics.
An important direction to extend the present work is to investigate solutions of the field equations (66$\land$(71). In particular, it would be interesting to see if a well-chosen form of the scalar function $\Lambda$ and vector field $V$ would allow, in a cosmological setup, to dynamically produce a behavior of the scale factor $a(t)$ compatible with the current universe accelerated expansion and/or with inflation.
In addition to that, it would also be necessary to study the behavior of compact objects (black holes, neutron stars) in our theory to see how these differ from the GR ones – and to which extent this remains compatible with modern observations.
@@ A.1 Symmetry
$\Phi:G\times M\to M:(g,x)\mapsto\Phi(g,x)\eqqcolon\Phi_{g}(x),$
@@
When $G$ is a Lie group, it is usually more convenient to study this notion of symmetry in terms of its infinitesimal generators. Let us denote by $\mathfrak{g}$ the Lie algebra of $G$ and by $X_{\xi}$ the generating vector field (on $M$), corresponding to a given $\xi\in\mathfrak{g}$ [^15] Technically, this means that $X_{\xi}$ is the vector field on $M$, whose integral curves are given by the flow of $\Phi$, i.e., at any point $p\in M$, the integral curve of $X_{\xi}$ through $p$ is $\mathscr{C}^{p}_{X_{\xi}}:t\mapsto\Phi(\exp(t\xi),p),$ where $\exp:\mathfrak{g}\to G$ is the exponential map.. If a tensor field $\boldsymbol{T}$ is invariant under the action of $G$, then
@@
Moreover, if $G$ is connected, relations (74) and (75) are actually, equivalent and one can then derive the conditions for a tensor field $\boldsymbol{T}$ to present a given symmetry via (75). It is the condition that we will use in the following.
Using this definition of symmetry, a metric-affine spacetime will be symmetric under a given group $G$ provided both the metric $\boldsymbol{g}$ and the distortion $\mathbf{L}$ are invariant under the action of $G$.
@@ A.2 Generators of cosmological symmetry
Cosmological symmetry refers to the idea that, following the cosmological principle, no observer has a prefered place in the universe, so that spacetime must be spatially homogeneous and isotropic. More precisely here, we then expect our setup to be invariant under both the “spatial” rotation group $SO(3)$ (providing isotropy) and a group representing spatial (quasi-)translations (providing spatial homogeneity).
One thus obtains 6 generators of cosmological symmetry $X_{i}$ $(i=1,\cdots,6)$. In local spherical coordinates $\left(x^{\mu}\right)=\left(t,r,\theta,\varphi\right)$ on $M$, $SO(3)$-generators are described as:
@@
as they satisfy the commutation relation of $\mathfrak{so}(3)$. Generators of (quasi-)translations are, [4],
@@
where $k\in\Set{-1,0,1}$ is the sign of the spatial curvature. When $k=0$ (flat spatial geometry), $X_{1}$, $X_{2}$ and $X_{3}$ correspond to the expression, in spherical coordinates, of the usual generators of spatial rotation around respectivelly the $x$-, $y$- and $z$-axis of an orthonormal Cartesian frame, while $X_{4}$, $X_{5}$ and $X_{6}$ correspond to the usual generators of spatial translation along respectivelly the $x$-, $y$- and $z$-axis.
@@ B.2.1 Variations with respect to the metric $\boldsymbol{g}$
When computing variation with respect to the metric, one has the usual relations, see e.g. [4],
@@
$\delta_{g}\accentset{\circ}{\Gamma}^{\rho}_{\mu\nu}$ $=\frac{1}{2}g^{\rho\alpha}\left(\accentset{\circ}{\nabla}_{\mu}\delta g_{\nu\alpha}-\accentset{\circ}{\nabla}_{\alpha}\delta g_{\mu\nu}+\accentset{\circ}{\nabla}_{\nu}\delta g_{\alpha\mu}\right),$ (92)